APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday September 23, 2007 (MJD 54367) ---=== OBSERVING TEAM ===--- Swing: Kaike Pan Night: Shannon Watters Steph Snedden ---=== OBSERVING PLAN ===--- SN Science! ---=== OBSERVING SUMMARY ===--- Clouds and fog. No science photons. Note that we let the spectrographs do an autofill during the night, and it behaved normally. ---=== OBSERVING LOG ===--- Afternoon: ---------- No problems to report. Night: ------ Clouds and high humidity persisted all night. We never went on-sky. We let the spectro autofill run tonight while we watched the fill. It shut off normally and stayed off. endNight rejoiced. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 6987 23:29Z 23:50Z 100 O 63.90 69.26 50 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n):n QA Procedures Done (y/n):n UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 46328, plate -9999 ------- 23:29 46328 0.0 bias ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 23:22Z 53 51 13 269 (W) 725 - cloudy 00:00Z 52 50 9 265 (W) 843 - ' 00:31Z 53 49 12 260 (W) 666 - ' 01:03Z 52 49 14 233 (SW) 882 - ' 01:35Z 52 49 13 256 (WSW) 857 - ' 02:07Z 52 49 11 267 (W) 768 - ' 02:39Z 52 49 10 270 (W) 779 - ' 03:11Z 51 49 14 258 (WSW) 842 - ' 03:44Z 51 48 11 260 (W) 688 - ' 04:16Z 50 48 12 241 (WSW) 632 - ' 04:48Z 50 48 10 241 (WSW) 671 - ' 05:20Z 50 48 10 263 (W) 620 - ' 05:52Z 50 48 14 257 (WSW) 587 - ' 06:24Z 50 48 16 221 (SW) 578 - ' 06:56Z 49 49 11 195 (SSW) 768 - ' 07:29Z 48 49 15 218 (SW) 594 - ' 08:01Z 48 49 13 233 (SW) 9460 - ' 08:33Z 48 49 18 225 (SW) 1777 - ' 09:05Z 49 50 17 249 (WSW) 6060 - ' 09:37Z 49 50 15 246 (WSW) 8605 - ' 10:09Z 49 50 18 269 (W) 8182 - ' ---=== TELESCOPE STATUS ===--- Never opened. Never ventilated Telescope stowed at 30 degrees with cart 4 mounted. Spectro autofill system turned off. The last fill was at 07:00Z, and should last well past the time the day crew arrives. Counterweights at 250 Spectro Dewar: 10:40Z 83 lb, 23 psi Imager Dewar : 10:40Z 226 lb, 22 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_38_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_36 SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.2050 -5.2130 Axial B -0.2154 -0.2135 Axial C -0.0015 -0.0016 Trans D 11.6350 11.6400 Lateral E 9.9568 9.9568 Lateral F 11.5316 11.5316 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: 10.9 Alt.: 30.342014 MIGS TONIGHT NOMINAL Axial A 0.7974 0.7973 Axial B 0.6046 0.6041 Axial C 0.5424 0.5425 Trans D 0.5997 0.5999 GALILS Commanded: 1798842. 1309017. 1643447. -9450. -4400. Actual: 1798628. 1309538. 1643409. -9472. -4376. SETMIR VALUES SecDesOrient: 1257.00 -60.00 -140.00 -190.00 200.12 SecOrient: 1256.93 -60.11 -140.21 -189.60 200.21 ---=== PROBLEMS IN DETAIL ===--- No unusual watcher messages - just the autolouver and adjVAXClock messages that we expect with this patched version of the TCC. No spectro focus was done this afternoon, since tonight was for Supernova.